OpenStudy (anonymous):

Im new to this site. x-y+z=-1 x+y+z= 5 x+y-z= 11 A) Nothing B) (5, -3, 3) C) (5, 3, -3) D) (-3, 5, 3)

hero (hero):

1. Create two sets of equations such that you are able to eliminate the same variable in both. In this case, we can create two sets of equations to eliminate the z variable by pairing equation one with equation three, then pairing equation 2 with equation three: x - y + z = -1 x + y - z = 11 x + y + z = 5 x + y - z = 11 Now adding both sets of equations together and this is what you get: 2x = 10 2x + 2y = 16 From there, you should be able to finish solving.

hero (hero):

You should easily be able to substitute 2x = 10 into the second equation. We already know what x equals. So solve this to get y: 10 + 2y = 16

hero (hero):

If you figure out x and figure out y, then figuring out z will be a no-brainer.

OpenStudy (anonymous):

ok ty, can u help me with another 1?

hero (hero):

Not so fast...you didn't confirm your answer to this one.

hero (hero):

You should let me know which answer choice you picked that way I know you understand it.

OpenStudy (anonymous):

oh i picked 5/3/-3

hero (hero):

Very good.

OpenStudy (anonymous):

u can u help with the next 1

hero (hero):

Is the next one similar to this question? If it is, then you should be able to figure it out on your own.

OpenStudy (anonymous):

its not.

hero (hero):

Okay, close this question, then post your next question as a new one.